# RoPINN: Region Optimized Physics-Informed Neural Networks

### v1

@@ 5 Conclusion
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\right]&=(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t}^{\prime})\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\right].\end{split}$ (30)
As for the second item on the right part, we also consider the upper bound of $\nabla_{\theta}\mathcal{L}(u_{\theta},\boldsymbol{x})$. However, different from point-wise optimization, $\boldsymbol{x}+\Omega$ could overlap with $\boldsymbol{x}^{\prime}+\Omega$. Thus, we can obtain the following estimation for the difference between the updated model parameters:
$\begin{split}&\mathbb{E}\left[G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\right]\\ &\leq\frac{|\Omega|-|\Omega_{r}|}{|\Omega|}\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\quad\quad\quad\text{(Separate analysis foroverlapped and non-overlap cases)}\\ &+\frac{|\Omega_{r}|}{|\Omega|}\mathbb{E}_{\Omega_{\mathrm{in}}}\left[\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &+\frac{|\Omega_{r}|}{|\Omega|}\mathbb{E}_{\Omega_{\mathrm{in}}}\left[\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\frac{|\Omega|-|\Omega_{r}|}{|\Omega|}\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]+\frac{|\Omega_{r}|}{|\Omega|}\mathbb{E}_{\Omega_{\mathrm{in}}}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L(1-\frac{|\Omega_{r}|}{|\Omega|}+\frac{|\Omega_{r}|-\frac{|\Omega_{r}|^{2}}{2^{(d+1)}}}{|\Omega|})\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L(1-\frac{|\Omega_{r}|^{2}}{2^{(d+1)}|\Omega|}),\end{split}$ (31)
where $\Omega_{\mathrm{in}}$ denotes the overlapped part between $\boldsymbol{x}+\Omega$ and $\boldsymbol{x}^{\prime}+\Omega$. In the above inequalities, the second inequality is based on the following derivation:
@@
$\begin{split}\left\|\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right\|&\leq\frac{|\Omega_{r}|}{|\Omega|}\frac{2}{\beta}\beta\|\theta_{t}-\theta_{t}^{\prime}\|\\ &\leq 2\|\theta_{t}-\theta_{t}^{\prime}\|.\end{split}$ (32)
@@
$\begin{split}\mathcal{E}_{\mathrm{gen}}\leq(1-\frac{|\Omega_{r}|^{2}}{2^{(d+1)}|\Omega|})\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}.\end{split}$ (33)
@@
Afterward, we will give a proof for the non-convex setting.
@@
For clarity, let $M=\frac{|\Omega_{r}|^{2}}{2^{(d+1)}|\Omega|}$. Then, we can rewrite the Eq. (25) as follows:
@@
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]&\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]\\ &+\frac{1}{|\mathcal{S}|}\left((1+\frac{M}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta t}(1-M)\right)\\ &\leq(1+\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|})\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|}(1-M)\\ &\leq\mathrm{exp}\left(\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|}\right)\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|}(1-M).\end{split}$ (34)

### v2

@@ 5 Conclusion
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\right]&=(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t}^{\prime})\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right],\end{split}$ (30)
where $G^{\mathrm{region}}_{\alpha,\boldsymbol{x}}(\theta)=\theta-\alpha\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(\theta,\boldsymbol{x})$. As for the second item on the right part, we also consider the upper bound of $\nabla_{\theta}\mathcal{L}_{r}^{\mathrm{region}}(u_{\theta},\boldsymbol{x})$. However, different from point-wise optimization, $\boldsymbol{x}+\Omega_{r}$ could overlap with $\boldsymbol{x}^{\prime}+\Omega_{r}$. For clarity, we define the overlapped area as $\Omega_{\mathrm{in}}$, whose size is larger than zero when $\boldsymbol{x}^{\prime}$ fall into the area centered at $\boldsymbol{x}$ with size $2^{(d+1)}|\Omega_{r}|$. Actually, due to the boundary of $\Omega$, we cannot always ensure $(\boldsymbol{x}^{\prime}+\Omega_{r})\subset\Omega$. Thus, for simplification, we assume that the domain $\Omega$ can be projected to a torus, where the out-of-domain samples will be re-included to $\Omega$.
Further, $\mathbb{E}_{\boldsymbol{x},\boldsymbol{x}^{\prime}\in\Omega,\ \mathrm{s.t.~{}}|\Omega_{\mathrm{in}}|=0}$ is simplified as $\mathbb{E}_{\Omega_{\mathrm{in}}=0}$ and $\mathbb{E}_{\boldsymbol{x},\boldsymbol{x}^{\prime}\in\Omega,\ \mathrm{s.t.~{}}|\Omega_{\mathrm{in}}|>0}$ is shorted as $\mathbb{E}_{\Omega_{\mathrm{in}}>0}$. And the operator $\mathbb{I}$ is defined as $\mathbb{I}(x)=\max\left(0,\min(1,x)\right)$ Thus, we can obtain the estimation for the difference between the updated model parameters through the following derivations:
@@
$\begin{split}&\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &=\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x^{\prime}}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &\leq\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{r}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L-\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L(1-\frac{|\Omega_{r}|}{|\Omega|}).\end{split}$ (31)
In the above inequalities, the third inequality is based on the following derivations. Firstly, denote $g=\|\left(\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\|$. According to Assumption 3.2, we have $g\leq\frac{|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}\beta\|\theta_{t}-\theta_{t}^{\prime}\|\leq\beta\|\theta_{t}-\theta_{t}^{\prime}\|$. Thus, the following inequality holds:
$\begin{split}&\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|^{2}\\ &=\|\theta_{t}-\theta_{t}^{\prime}\|^{2}-2\alpha_{t}g^{\sf T}(\theta_{t}-\theta_{t}^{\prime})+\alpha_{t}^{2}g^{2}\\ &\leq\|\theta_{t}-\theta_{t}^{\prime}\|^{2}-2\frac{\alpha_{t}}{\beta}g^{2}+\alpha_{t}^{2}g^{2}\quad\quad\quad\quad\quad\quad\quad\text{({Convexity} and $g\leq\beta\|\theta_{t}-\theta_{t}^{\prime}\|$)}\\ &\leq\|\theta_{t}-\theta_{t}^{\prime}\|^{2}.\quad\quad\quad\quad\ \ \quad\quad\quad\quad\quad\quad\ \quad\quad\quad\ \text{($\alpha_{t}\leq\frac{2}{\beta}$)}\end{split}$ (32)
@@
$\begin{split}\mathcal{E}_{\mathrm{gen}}\leq(1-\frac{|\Omega_{r}|}{|\Omega|})\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}.\end{split}$ (33)
For the more general case, we no longer assume that the domain $\Omega$ can be projected to a torus, resulting in a non-symmetric scenario when $|\Omega_{\mathrm{in}}|>0$. This asymmetry arises due to the presence of boundaries, since points that could potentially intersect with the set $\boldsymbol{x}+\Omega_{r}$ may be truncated by the boundary. Specifically, we consider $\Omega=[0,l]^{(d+1)}$ and $\Omega_{r}=[0,r]^{(d+1)}$. The concrete probability $\mathbb{P}(|\Omega_{\mathrm{in}}|>0)$ and the expectation $\mathbb{E}_{|\Omega_{\mathrm{in}}|>0}(|\Omega_{\mathrm{in}}|)$ can be calculated as follows:
$\begin{split}\mathbb{P}(|\Omega_{\mathrm{in}}|>0)=(\frac{r(2l-3r)}{(l-r)^{2}})^{(d+1)},\ \mathbb{E}_{|\Omega_{\mathrm{in}}|>0}(|\Omega_{\mathrm{in}}|)=(\frac{r^{2}(3l-4r)}{3(l-r)^{2}})^{(d+1)}.\end{split}$ (34)
Thus, the general case of Eq. 31 can be reformulated using the identities above. We assume $\frac{r}{l}<0.5$, as when $\frac{r}{l}\geq 0.5$, it follows that $\mathbb{P}(|\Omega_{\mathrm{in}}|>0)=1$. Specifically, Eq. (31) can be rewrite as:
$\begin{split}&\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &=\mathbb{P}(|\Omega_{\mathrm{in}}|=0)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]+\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq\mathbb{P}(|\Omega_{\mathrm{in}}|=0)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\\ &+\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x^{\prime}}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &\leq\mathbb{P}(|\Omega_{\mathrm{in}}|=0)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]+\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &+\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &\leq\mathbb{P}(|\Omega_{\mathrm{in}}|=0)\mathbb{E}_{\Omega_{r}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]+\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L-\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L\left[1-\mathbb{P}(|\Omega_{\mathrm{in}}|>0)\frac{\mathbb{E}_{|\Omega_{\mathrm{in}}|>0}(|\Omega_{\mathrm{in}}|)}{|\Omega_{r}|}\right]\\ &=\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L\left[1-(\frac{r^{2}(2l-3r)(3l-4r)}{3(l-r)^{4}})^{(d+1)}\right].\end{split}$ (35)
$\begin{split}\mathcal{E}_{\mathrm{gen}}\leq\left[1-(\frac{r^{2}(2l-3r)(3l-4r)}{3(l-r)^{4}})^{(d+1)}\right]\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}.\end{split}$ (36)
Although the specific forms differ and the general case is much more complex, these two inequalities both share the same intuitive meaning: region optimization benefits from the overlap between $\boldsymbol{x}+\Omega_{r}$ and $\boldsymbol{x}^{\prime}+\Omega_{r}$. Moreover, within a certain range, the benefit increases as the value of $r$ becomes larger.
Thus, to keep the bound simple and easy to understand, the main text theorems are under the assumption that $\Omega$ can be projected to a torus. Otherwise, $\frac{|\Omega_{r}|}{|\Omega|}$ should be replaced by $(\frac{r^{2}(2l-3r)(3l-4r)}{3(l-r)^{4}})^{(d+1)}$.
@@
Next, we will prove the non-convex setting. Similarly, we assume $\Omega$ can be projected to a torus. Otherwise, the $\frac{|\Omega_{r}|}{|\Omega|}$ in the final bound should be replaced by $(\frac{r^{2}(2l-3r)(3l-4r)}{3(l-r)^{4}})^{(d+1)}$.
@@
For clarity, let $M=\frac{|\Omega_{r}|}{|\Omega|}$. Then, we can rewrite the Eq. (25) as follows.
@@
$\begin{split}\mathbb{E}\left[\|\theta_{t+1}-\theta_{t+1}^{\prime}\||\delta_{t_{0}}=0\right]&\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq(1-\frac{1}{|\mathcal{S}|})(1+\frac{1}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{1}{|\mathcal{S}|}\left((1+\frac{M}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta t}(1-M)\right)\\ &\leq(1+\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|})\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|}(1-M)\\ &\leq\mathrm{exp}\left(\frac{1}{t}-\frac{1-M}{t|\mathcal{S}|}\right)\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|}(1-M),\end{split}$ (37)
where the second inequality is based on the following derivations:
$\begin{split}&\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &=\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]\\ &\leq\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|\theta_{t}-\theta_{t}^{\prime}-\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{r}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &\leq\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|\theta_{t}-\theta_{t}^{\prime}\right\|+\left\|\left(\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}-\alpha_{t}\frac{1}{|\Omega_{r}|}\int_{\Omega_{\mathrm{in}}}\nabla_{\theta}\mathcal{L}({u_{\theta_{t}^{\prime}}},\boldsymbol{x}+\boldsymbol{\xi})\mathrm{d}\boldsymbol{\xi}\right)\right\|\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{I}\left(\frac{|\Omega|-2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{r}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]\\ &+\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\frac{|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}\alpha_{t}\beta\|\theta_{t}-\theta_{t}^{\prime}\|+\frac{|\Omega_{r}|-|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{\alpha_{t}\beta|\Omega_{r}|}{|\Omega|}\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L-\mathbb{I}\left(\frac{2^{(d+1)}|\Omega_{r}|}{|\Omega|}\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{|\Omega_{\mathrm{in}}|}{|\Omega_{r}|}2\alpha_{t}L\right]\\ &\leq(1+\frac{\alpha_{t}\beta|\Omega_{r}|}{|\Omega|})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L(1-\frac{|\Omega_{r}|}{|\Omega|})\\ &=(1+\frac{M}{t})\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta t}(1-M).\end{split}$ (38)
Notably, $\mathbb{E}\left[\|G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}}(\theta_{t})-G^{\mathrm{region}}_{\alpha_{t},\boldsymbol{x}^{\prime}}(\theta_{t}^{\prime})\|\right]$ has an obvious upper bound, i.e. $\left(\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+\frac{2L}{\beta t}\right)$. And only when $\mathbb{E}(\delta_{t})\leq\frac{2L}{\beta}$, the bound derived by Eq. (38) is tighter. Furthermore, the condition that $\mathbb{E}(\delta_{t})\leq\frac{2L}{\beta}$ can be easily satisfied at the beginning several iterations since $\mathbb{E}(\delta_{0})=0$.